# Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Multi-modal Text Recognition

### v3

@@ 3.2 Fusing text-recognition models with LLM
To evaluate the efficacy of our algorithm, we implemented GFD for ASR and OCR tasks. This is achieved by fusing pre-trained text-recognition models with LLMs to enhance recognition capability. In such scenarios, the proposed tokens of ASR and OCR model are conditioned on a modal signal $\mathcal{Z}^{\texttt{TR}}$ while that of the LLM is not. Therefore, the same next-byte probability estimation from the LLM and text recognition model might stem from different token choices. Although the difference would converge eventually as the sequence rolls out, given limited beams in the beam search process, the potential overestimation of implausible words due to scores contributed by LLM spikes the probabilities of common word-leading bytes such as ’t’ and ’r’, and produces locally unstable rankings that may result in false eliminations of correct beams in decoding steps. In response to this challenge, we have adopted text recognition models (ASR and OCR) to propose the next byte/token, while employing LLMs to provide delayed corrective feedback on past tokens. The fusion decoding methodology we have adopted, derived from Equation (5), is delineated in the following formula:
$\{B_{1},..,B_{t}\}^{*}=\underset{\{B_{1},..,B_{t}\}}{\mathrm{arg\,max}}\bigl{[}(1-r)\cdot\log\mathbb{P}^{\texttt{TR}}(\{B_{1},..,B_{t}\},\mathcal{Z}^{\texttt{TR}})$ (8)
@@
where $\{B_{1},\ldots,B_{t}\}^{*}$ represents the optimal sequence of bytes up to and including position $t$, $p^{\texttt{TR}}$ and $p^{\texttt{LM}}$ are the probability distributions of the text-recognition models and the language models, respectively, $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ are the contextual information for the text-recognition and language models, respectively, $r$ is a weighting factor that balances the influence of the text-recognition models and the language models, and $k$ depends on the length of the byte/token proposal. As shown in Figure 1, $\mathcal{Z}^{\texttt{TR}}$ can be either an audio or a text image, and $\mathcal{Z}^{\texttt{LM}}$ can be a prompt.
4 Experiment
@@ 4.1 Experimental setup
We evaluate the application of GFD to ASR and OCR tasks. For ASR task, we benchmark datasets in English, Chinese, and Cantonese. The deliberate choice of Chinese and Cantonese is due to that the languages are homophonous and tonal, challenging for ASR systems. This complexity is corroborated by the Fleurs experiment in Whisper [40], where the Chinese word error rate is well above the regressed word error rate in comparison to evaluations in other language at the same amount of pre-training data. Of all tested languages in Whisper, it is the only large-scale language (> 10k hours audio) with such a phenomenon. For OCR, we benchmark image dataset containing long sequence of text as we hypothesize that LLM provide semantic information to an OCR model of recognizing long text sequences.
@@
In terms of the evaluated models, we choose Whisper-large-v2 as the ASR model for greedy and beam search. For our proposed GFD, we use Mistral [26] for English datasets and Breeze [22] for Chinese and Cantonese datasets. We term the GFD fused model with Whisper and Mistral as GFD-ASR-EN; that with Whisper and Mistral as GFD-ASR-ZH. In addition, we benchmark GER, based on task-activating prompting method [5, 48], with Mistral-Instruct and Breeze-Instruct models without finetuning. We include two oracle word error rates following previous work[24], where the N-Best Oracle $o_{nb}$ denotes the error rate calculated with the best candidate in the N-best list, and the Compositional Oracle $o_{cp}$ is the best achievable word error rate using all tokens N-best the list. As to OCR, we utilize TrOCR [32], a pre-trained OCR model, and denote the fused model GFD-OCR.
@@
We benchmark models on various datasets, including Librispeech [21], Medical [18], ATCO2 [45], Fleurs [12], NTUML2021[49], and FormosaSpeech  [^2] https://huggingface.co/datasets/MediaTek-Research/formosaspeech for ASR; NAF [13] for OCR. Librispeech is a collection of corpus from audiobooks with subsets Clean and Other. The Librispeech-Noise is its noised variant, featuring various signal-to-noise ratios, ideally for testing ASR system’s robustness to noise. Medical dataset contains 8.5 hours of medical conversations with associated symptom tags to each audio-text pairs. ATCO2 contains audios of air traffic control communication and accompanied meta-information of airports. Fleurs is a multilingual speech corpus. We deliberately choose Cantonese subset for evaluation as the language is homophonous and tonal. NTUML2021 corpus consists of lecture recordings from the “Machine Learning” course at National Taiwan University in 2021, with corresponding transcriptions and English translations labeled by over 20 bilingual native Chinese speakers. FormosaSpeech corpus includes Chinese recordings of Taiwanese accents amassing up to 6.4 hours of audio-text pairs. NAF consists of images from U.S national archives with labelled bounding boxes and annotations, ideal for evaluating OCR performance.
For all ASR experiments, we report word error rate (WER) for evaluations on English datasets, and mix error rate (MER) for that on Chinese datasets. For OCR experiments, we report character error rate (CER) and exact match (EM). For exact match, we require that the recognized texts fully matches the ground truth. All GFD fused models can be run on a single A6000 GPU. We fix $r=0.2$ in Equation 8 for all our experiments; number of beams equaling 5 and 10 for ASR and OCR experiments, respectively.
4.2 Conventional speech recognition
We first verify the efficacy of GFD in conventional speech recognition setting and report results in Table 1. We notice that beam search improves consistently upon greedy search, and use beam search as the baseline. We found that GFD moderately improves on the Medical and Fleurs-HK dataset. In the more challenging NTUML2021, we obtained a state-of-the-art 8.83 mixed error rate, surpassing prior work[49] and even the oracle N-Best score. This improvement is mostly preserved even when we narrow the target space to only consider English terms in the dataset (Table 2), making GFD-ASR an elegant solution for code-switched ASR. Examples in the Appendix show that improvement can be attributed to the LLM correcting grammatical mistakes and domain terminologies. Performance across Librispeech reveals that GFD improves most on moderate noise conditions but diminishes when the noise is too large $(S/R=-5)$.
In addition, we do not find improvements in the GER setting using non-GER-specific instruct models, consistent with previous work[5]. As we already patched malformatted outputs with the best beam result, this increase of error rate is mostly attributed to LLM hallucinations, including incorrect deletions. We therefore posit that GFD is more robust for incorporating off-the-shelf LLMs on the ASR task than GER.
Table 1: Results on conventional speech recognition.
Whisper GER GFD-ASR-EN Oracle
EN dataset Greedy 5-beams (Baseline) $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
Librispeech-Clean 2.30 2.28 3.21 2.29 1.91 1.63
Librispeech-Other 5.23 4.97 6.57 4.99 4.20 3.43
Medical 7.30 7.22 7.49 6.74 6.17 5.05
Librispeech-Noise ($S/R=10$) 3.50 3.12 4.05 3.02 2.32 1.94
Librispeech-Noise ($S/R=5$) 5.67 5.25 6.00 4.96 4.10 3.28
Librispeech-Noise ($S/R=0$) 15.23 13.54 14.44 13.27 11.66 9.04
Librispeech-Noise ($S/R=-5$) 49.09 47.05 47.35 46.98 43.62 33.19
Whisper GER GFD-ASR-ZH Oracle
ZH or HK dataset Greedy 5-beams (Baseline) $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
Fleurs-HK (Cantonese) 7.49 6.88 7.33 6.23 5.58 4.58
NTUML2021 11.11 9.97 9.87 8.83 8.88 4.54
NTUML2021 FormosaSpeech
Full English
Whisper (short) 9.97 15.21 -
SICL (short) [49] 9.5 - -
GFD-ASR-ZH (short) 8.83 13.51 -
Whisper (long) 8.40 11.56 27.41
GFD-ASR-ZH (long) 8.18 11.43 26.72
Table 2: Comparison results on long-form and short-form.
4.3 Long-form speech recognition
As LLMs are capable in attending to long context of sequence and possess multilingual capabilities, we posit that GFD-ASR-ZH can improve on long-form audio transcription. We therefore evaluate the long-form transcription performance on NTUML2021 and FormosaSpeech. In constrast to the short-form evaluations in Table 1 in NTUML2021, we prepend all previous transcription from contiguous clips in the same lecture as prompts for both ASR and LLM to realize long form transcriptions, but truncate it when necessary for whisper due to context length limitations. In the setting of evaluations on contiguous short clips, GFD-ASR-ZH (long) consistently outperforms Whisper (long) in both NTUML2021 and FormosaSpeech, demonstrating that the long-context capability of LLM can be effectively utilized through GFD.
4.4 Context-aware speech recognition
Table 3: Results on context-aware ASR task. We set num beams=5 across all our experiments.
@@ 4.4 Context-aware speech recognition
Whisper No - 47.70 66.44 2.28 4.97 27.42 7.22
@@
In context-aware speech recognition, we aim to investigate the capability of the fusion model in leveraging contextual information, which is highly desired in real-life contexts, where the incoming speech is expected to be bounded by domain, contain certain specific words, or follow a certain transcribing format. We utilize prompting with ASR model and the LLM model (in the GFD setting) to condition on these external hints, in three following settings.
@@
Domain and Rare Word Prompting: We tested domain-conditioned ASR on the Medical dataset, where the symptom tags are provided along with the speech content. Results show that our GFD method with LLM prompting improves on GFD without prompting, showing prompt sensitiveness of the LLM in the GFD system. However, we did not find further improvement upon whisper prompting, compared to double-prompted GFD. For verifying rare word prompting capabilities, we used the augmented Librispeech and FormosaSpeech dataset, where a target rare word is mixed with 100 other distractors for each data point [28]. For the FormosaSpeech Dataset, we created the rare words with ChatGPT, and generate distractors with a similar approach using the training set. At this scale of distractors, it is unrealistic to prompt on Whisper, as the remaining context length is often insufficient for ASR decoding. By prompting on the LLM using GFD, we demonstrated up to 6% WERR over non-prompting methods on the Librispeech dataset, and 1.6% WERR on the FormosaSpeech Dataset.
@@
Instruction Aware Prompting: We evaluated formatted speech recognition on the ATCO2 dataset, a dataset on air traffic control communications, which has strict regulations on call signs and transcribe formats. By incorporating a LLM, we are able to prompt it with over 4000 words of guidelines from an enitre instruction manual [^3] https://www.faa.gov/air_traffic/publications/atpubs/aim_html/chap4_section_2.html. a possibility not present with Whisper. For Whisper prompting, we include all special call signs and three example sentences extracted from the manual, where we leave details in the Appendix. Results in Table 3shows that GFD with ASR and LLM prompting obtains best results with a WERR of 17.7% compared with Whisper in the normalized (ATCO2-Norm) setting, even outperforming Clairaudience[35], a fine-tuned prompt conditioning model. We also reported scores without word conversion normalization (ATCO2-Raw), from observations that standard normalization can excessively correct errors that conflicts with the transcribing guidelines. In this setting, our gains are even larger, showing instruction following abilites of the GFD system.
@@ 4.5 Optical Character Recognition
Table 4: Evaluation on OCR task
@@
We demonstrate that GFD can be applied to fusing end-to-end OCR model with LLMs, using TrOCR and Mistral [32, 26]. We experiment the fusion model on a dataset extracted from National Archive Forms dataset [13], denoted NAF-Long. The NAF-Long dataset consists of images with text sequences having character lengths between 100 and 200, equivalent to around 25 to 50 text tokens. In Table 4, we show that fusing LLM to the OCR model significantly improves the OCR results by 16.7 % in character error rate reduction and 38.07 % in exact match improvement.
5 Discussion
@@
Further comparisons with GER. The GFD algorithm relates to GER in that both algorithms perform the selection of output sequences with beam decoding. However, they differ in compute execution and the diversity of sample sequences. First, in GFD, the LLM works in parallel with the ASR decoder, and thus the computation of per step inference can be executed asynchronously with a load bounded by $\mathcal{O}(Z)+\mathcal{O}(k\cdot\max(S_{ASR},{S_{LLM}}))$, where $Z$ denotes the size of speech encoding, $k$ is the beam size, and $S_{ASR}$ and $S_{LLM}$ are ASR and LLM decoding lengths, respectively. The LLM decoding complexity expression matches that of Section 3.2, treating the vocabulary size as a constant in this dicussion. In contrast, GER operates sequentially, requiring the completion of beam decoding with ASR prior to a correction with LLM. This means the execution time of GER is bounded by $\mathcal{O}(Z)+\mathcal{O}((k+1)\cdot S_{LLM})+\mathcal{O}(k\cdot S_{ASR})$, where the additional $O(S_{LLM})$ comes from LLM decoding. Secondly, in GFD, the searched token space attended by the LLM is at least n_beams * sequence length, whereas in GER, the top-k cutoff of the ASR step does not promote diversity between the candidates, which limits the LLM search space. Aside from these differences, GFD and GER are methodologically orthogonal, allowing for their combination in the pursuit of further improvement.

### v4

@@ 3.2 Fusing text-recognition models with LLM
To evaluate the efficacy of our algorithm, we implemented GFD for ASR and OCR tasks. This is achieved by fusing pre-trained text-recognition models with LLMs to enhance recognition capability. Essentially, the text recognition models (ASR and OCR) propose sequences for the LLMs to provide scoring feedback. To limit the number of proposals scored by the LLM for reasonable time complexity, we introduce a delayed corrective feedback loop to coordinate the two models, characterized by a dynamic shifting value $k$. Based on Equation (5), the fusion decoding methodology used in our experiments is given by the following formula:
$\{B_{1},..,B_{t}\}^{*}$ (8)
$=\underset{\{B_{1},..,B_{t}\}}{\mathrm{arg\,max}}\bigl{[}$ $(1-r)\cdot\log\mathbb{P}^{\texttt{TR}}(\{B_{1},..,B_{t}\},\mathcal{Z}^{\texttt{TR}})$
@@
where $\{B_{1},\ldots,B_{t}\}^{*}$ represents the optimal sequence of bytes up to and including position $t$, $p^{\texttt{TR}}$ and $p^{\texttt{LM}}$ are the probability distributions of the text-recognition models and the language models, respectively, $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ are the contextual information for the text-recognition and language models, respectively, $r$ is a weighting factor that balances the influence of the text-recognition models and the language models, which is determined via grid search on a small scale experiment (Appendix  A.2). $k$ is optimally selected to be equal to the length of the last token of the proposal from $\mathcal{M}^{\texttt{TR}}$ (Appendix  A.3). We note that even with a shifting value $k$, the necessity of equation 7 still holds.
4 Experiments
Whisper Re-ranking GER GFD-ASR-EN Oracle
EN dataset Greedy 5-beams $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
Librispeech-Clean 2.30 2.28 2.20 2.41 2.29 1.91 1.63
Librispeech-Other 5.23 4.97 4.86 5.30 4.99 4.20 3.43
Medical 7.30 7.22 7.22 7.49 6.74 6.17 5.05
Librispeech-Noise ($S/R=10$) 3.50 3.12 3.27 3.14 3.02 2.32 1.94
Librispeech-Noise ($S/R=5$) 5.67 5.25 5.40 5.38 4.96 4.10 3.28
Librispeech-Noise ($S/R=0$) 15.23 13.54 13.69 13.70 13.27 11.66 9.04
Librispeech-Noise ($S/R=-5$) 49.09 47.05 47.04 47.35 46.98 43.62 33.19
Whisper Re-ranking GER GFD-ASR-ZH Oracle
ZH or HK dataset Greedy 5-beams $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
Fleurs-HK (Cantonese) 7.49 6.88 7.00 7.33 6.23 5.58 4.58
NTUML2021 11.11 9.97 9.68 9.87 8.83 8.88 4.54
Table 1: Performance for short-form speech recognition. The table presents Word Error Rate (WER) for EN and Mixed Error Rate (MER) for ZH/HK. Bold values indicate the best performance excluding the Oracle column. Re-ranking and GER utilize LLMs with 5-beam outputs from Whisper, whereas GFD integrates LLMs during the decoding process.
Librispeech-Noise (S/R=0)
Whisper-5beams 13.54
RobustGER Hu et al. (2024) 13.20
GFD 13.27
RobustGER+GFD 13.03
Table 2: Comparisons on GER and GFD. GFD and GER improvements similarly on Whisper, where ensembling of both approaches performs best.
@@ 4.1 Experimental setup
We evaluate the application of GFD to ASR and OCR tasks. For ASR task, we benchmark datasets in English, Taiwanese Mandarin, and Cantonese. The deliberate selection of Taiwanese Mandarin and Cantonese is due to their homophonic and tonal characteristics, which reveal robustness shortcomings of ASR systems. This complexity is corroborated by the Fleurs experiment in Whisper Radford et al. (2023), where the Chinese word error rate is well above the regressed word error rate in comparison to evaluations in other language at the same amount of pre-training data. Of all tested languages in Whisper, it is the only large-scale language (more than 10k hours audio) with such a phenomenon. For OCR, we benchmark image dataset containing long sequence of text as we hypothesize that LLM provide semantic information to an OCR model of recognizing long text sequences.
@@
For the evaluated models, we selected Whisper-large-v2 as the ASR model for both greedy and beam search methods. In our proposed GFD approach, we utilized Mistral Jiang et al. (2023) as the language model for English datasets, referring to this configuration as GFD-ASR-EN. For Chinese and Cantonese datasets, we integrated Breeze Hsu et al. (2024) and designated this setup as GFD-ASR-ZH. In addition, we benchmark GER, based on task-activating prompting method Chen et al. (2024a); Yang et al. (2023b), with Instruction-tuned models including Mistral-Instruct. We include two oracle word error rates following previous work Hu et al. (2024), where the N-Best Oracle $o_{nb}$ denotes the error rate calculated with the best candidate in the N-Best list, and the Compositional Oracle $o_{cp}$ is the best achievable word error rate using all tokens N-Best the list. As to OCR tasks, we utilize TrOCR Li et al. (2023a) with Mistral and denote the fused model GFD-OCR-EN.
@@
We benchmark the models on a wide variety of datasets, including Librispeech Han et al. (2019), Medical Figure Eight Inc. (2019), ATCO2 Szöke et al. (2021), Fleurs Conneau et al. (2023), NTUML2021Yang et al. (2023c), and FormosaSpeech for ASR; NAF Davis et al. (2019) for OCR. Librispeech is a collection of corpus from audiobooks with subsets "Clean" and "Other". Librispeech-Noise is a noised variant of the original LibriSpeech dataset with different signal-to-noise ratios, which is ideal for testing ASR systems’ robustness to noise. Medical dataset contains 8.5 hours of medical conversations with associated symptom tags to each audio-text pairs. ATCO2 contains audios of air traffic control communication and accompanied meta-information of airports. Fleurs is a multilingual speech corpus. We deliberately choose Cantonese subset for evaluation as the language is homophonous and tonal. NTUML2021 corpus consists of lecture recordings from the “Machine Learning” course at National Taiwan University in 2021, with corresponding transcriptions and English translations labeled by over 20 bilingual native Chinese speakers. FormosaSpeech corpus includes Chinese recordings of Taiwanese accents amassing up to 6.4 hours of audio-text pairs. NAF consists of images from U.S national archives with labelled bounding boxes and annotations, ideal for evaluating OCR performance. For all ASR experiments, we report Word Error Rate (WER) for evaluations on English datasets and Mixed Error Rate (MER) for those on Chinese datasets. For OCR experiments, we report Character Error Rate (CER) and Exact Match (EM).
@@
RobustGER, UADF Chen et al. (2024b) No Yes >100 >100 >100 >100 >100 >100
Clairaudience Liao et al. (2023) Yes - 28.77 - - - - 6.54
RobustGER Hu et al. (2024) Yes No 34.77 50.58 - - - -
Whisper No - 47.70 66.44 2.28 4.97 22.33 7.22
@@
Table 3: Results on instruction-aware ASR task. All experiments are done with a beam size of 5. These experiments is conditioned on a given prompt containing either domain tags (on Medical), rare words (on Librispeech and FormosaSpeech), and complex transcription guidelines (on ATCO2). Error rates of over 100 is reported with GER-based methods (Hu et al., 2024; Chen et al., 2024b), as they catastrophically fail on prompt conditioning due to their inability to process instruction prompts beyond the GER prompt.
4.2 Short-form speech recognition
We verify the efficacy of GFD in speech recognition setting and report results in Table 1. First, we notice that for decoding results without incoporating an LLM, beam search improves consistently upon greedy search. Using beam search as the baseline, we observe that GFD moderately improves on the Medical and Fleurs-HK dataset. In the more challenging NTUML2021, we obtained a 8.83 mixed error rate, surpassing even the oracle N-Best score. Upon examining the benchmarked samples, we attribute the observed improvements to the LLM’s ability to correct English grammatical mistakes and domain-specific terminology within the code-switching context. As such, GFD serves as an elegant solution that facilitates the success of code-switched ASR systems. Performance across Librispeech demonstrates that GFD offers the most significant enhancement under moderate noise conditions but diminishes when the noise level is too high $(S/R=-5)$.
In contrast, we do not find improvements in the GER setting using general instruct models, consistent with previous work Chen et al. (2024a). The increased error rate is primarily attributed to LLM hallucinations, including incorrect deletions. Therefore, we posit that GFD is more robust than GER for incorporating off-the-shelf LLMs in the ASR task. As demonstrated in Table 2, GFD achieves similar improvements when compared to a specialized GER model, RobustGER Hu et al. (2024). Analyzing the outputs reveals that the corrected errors are orthogonal: while GER is specifically instructed to adhere to the words in the N-best list, GFD can select words from an exponential search space with intermediate interrogation. The combination of RobustGER and GFD yields the best results.
4.3 Long-form speech recognition
Whisper (5-beams) GFD-ASR-ZH
NTUML2021 (long-form) 8.40 8.18
FormosaSpeech 22.33 20.59
Table 4: Performance for long-form speech recognition. For NTUML2021, we concatenate all contiguous clips to reconstruct the original lecture for long-form evaluation.
LLM’s capability to attend to long sequences, makes it an appealing candidate on long-form audio speech recognition. Therefore, we evaluate long-form transcription performance on NTUML2021 and FormosaSpeech. For NTUML2021, we curate the long dataset by concatenating all contiguous clips to reconstruct the original lecture for long-form evaluation. To properly contrast the short-range modeling in Table 1, we prepend all historical transcriptions as prompts for ASR and LLM to realize long-form transcriptions, truncating them when necessary for Whisper due to context length limitations. In this setting, GFD-ASR-ZH consistently outperforms Whisper in both NTUML2021 and FormosaSpeech, demonstrating that the long-context capability of LLM can be effectively utilized through GFD (Table 4).
4.4 Instruction-aware speech recognition
In instruction-aware speech recognition, we explore GFD’s ability to leverage contextual information, which is crucial in real-life scenarios where speech may be domain-specific, contain rare or critical terms, or require adherence to complex transcription guidelines. We employ prompting with ASR and LLM models along with GFD to incorporate these external cues across the respective three settings.
@@ 4.4 Instruction-aware speech recognition
Domain tag and rare word prompting. We tested domain-conditioned ASR on the Medical dataset, where the symptom tags are provided along with the speech content. Results show that our GFD method with LLM prompting improves on GFD without prompting, showing prompt sensitiveness of the LLM in the GFD system. However, we did not find further improvement upon whisper prompting, compared to double-prompted GFD. For verifying rare word prompting capabilities, we used the augmented Librispeech and FormosaSpeech dataset, where a target rare word is mixed with 100 other distractors for each data point Le et al. (2021). For the FormosaSpeech Dataset, we created the rare words with ChatGPT, and generate distractors with a similar approach using the training set. At this scale of distractors, it is unrealistic to prompt on Whisper, as the remaining context length is often insufficient for ASR decoding. By prompting on the LLM using GFD, we demonstrated up to 7% WERR over non-prompting methods on the Librispeech dataset, and 1.6% WERR on the FormosaSpeech Dataset (Table 3).
@@
Prompting with Instructions. We evaluated formatted speech recognition on the ATCO2 dataset, a dataset on air traffic control communications, which has strict regulations on call signs and transcribe formats. By incorporating an LLM, we are able to prompt it with over 4000 words of guidelines from an entire instruction manual [^2] https://www.faa.gov/air_traffic/publications/atpubs/aim_html/chap4_section_2.html, a possibility not present with Whisper. For Whisper prompting, we include all special call signs and three example sentences extracted from the manual (Appendix A.2). Results in Table 3 show that GFD with ASR and LLM prompting obtains best results with a WERR of 17.7% compared with Whisper in the normalized (ATCO2-Norm) setting, even outperforming Clairaudience Liao et al. (2023), a fine-tuned prompt conditioning model. GER-based methods Hu et al. (2024); Chen et al. (2024b) also fall short in this category, due to their inability to process instruction prompts beyond the GER prompt. Transcription results completely diverge from the spoken content, causing meaningless error rates exceeding 100%. We also reported scores without word conversion normalization (ATCO2-Raw), from observations that standard normalization, such as converting arabic to written numerals, can excessively correct errors that conflicts with the transcribing guidelines. In this setting, our improvements are even more pronounced, further demonstrating the instruction-following capabilities of the GFD system.
@@ 4.5 Optical Character Recognition
We use the OCR task as an example to demonstrate that GFD is applicable to auto-regressive scenarios beyond ASR. In Table 5, we show that fusing the Mistral LLM to the TrOCR model significantly improves the OCR results on the National Archive Forms dataset by 16.7 % in character error rate reduction and 38.07 % in exact match improvement.
@@
Table 5: Evaluation on NAF-Long (an OCR task).
5 Analysis
5.1 Further comparisons with GER
@@
The GFD algorithm relates to GER in that both algorithms perform the selection of output sequences with beam decoding. However, they differ in compute execution and the diversity of sample sequences. First, in GFD, the LLM works in parallel with the ASR decoder, and thus the computation of per step inference can be executed asynchronously with a load bounded by $\mathcal{O}(Z)+\mathcal{O}(k\cdot\max(S_{ASR},{S_{LLM}}))$, where $Z$ denotes the size of speech encoding, $k$ is the beam size, and $S_{ASR}$ and $S_{LLM}$ are ASR and LLM decoding costs of a single token, respectively. The LLM decoding complexity expression matches that of Section 3.2, treating the vocabulary size as a constant in this dicussion. In contrast, GER operates sequentially, requiring the completion of beam decoding with ASR prior to a correction with LLM. This means the execution time of GER is bounded by $\mathcal{O}(Z)+\mathcal{O}((k+1)\cdot S_{LLM})+\mathcal{O}(k\cdot S_{ASR})$, where the additional $O(S_{LLM})$ comes from LLM decoding. Secondly, in GFD, the searched token space attended by the LLM is at least n_beams times sequence length, whereas in GER, the top-k cutoff of the ASR step does not promote diversity between the candidates, which limits the LLM search space. Aside from these differences, GFD and GER are methodologically orthogonal, allowing for their combination in the pursuit of further improvement.
5.2 Further comparisons with Other Late Fusion Approaches
Concurrently with our work, step-by-step synchronous late fusion methods that focus on resolving tokenization mismatch have been explored Mittal et al. (2024); Hori et al. (2025). However, these approaches impose constraints on the scoring process, limiting it to specific decoding configurations. For example, in SALSA Mittal et al. (2024), the LLM sequence is only rescored if it is ending in a UTF-8 character, a criterion designed for non-ascii languages such as Mandarin and Hindi. Similarly, Delayed Fusion Hori et al. (2025) discards trailing partial tokens from the ASR that do not form complete words during LLM rescoring, which is not easily extensible to languages without explicit word boundaries, such as Mandarin. In contrast, our method approaches the problem from the byte sequence level, enabling a more general and flexible rescoring process that supports arbitrary input sequences. From an information-theoretic perspective, our approach preserves the maximum available information at each decoding step, avoiding sequence cutoffs or step skipping. Empirical results in Table 6 support our hypothesis, with GFD outperforming other settings across English and Mandarin. As further detailed in Appendix A.1, the computational overhead introduced by branching through GFD is negligible compared to the cost of a single forward pass of the large language model.
Method Libri.-Noisy-5 (en) Formosa-Sp. (zh)
GFD (Ours) 4.96 20.59
No Branching 4.98 21.29
+ Word Cutoff 5.18 22.33x
+ Char. Cutoff 4.98y 21.50
Table 6: Word error rates of GFD compared with alternative rescoring methods across English and Mandarin. After removing branching, additional constraints can be applied—such as discarding trailing partial tokens, similar to Hori et al. (2025), or removing trailing bytes that do not form complete UTF-8 characters, similar to Mittal et al. (2024).xThis setting reduces to no rescoring due to the absence of explicit word boundaries. yThis setting is identical to no branching in ASCII languages.
5.3 Error Analysis